# Select, then Balance: A Plug-and-Play Framework for Exogenous-Aware Spatio-Temporal Forecasting

### v1

@@ 3.3.1. Siamese ST Encoders
where and encapsulate the high-level spatio-temporal features derived from past and future exogenous contexts, respectively, preparing them for adaptive fusion.
3.3.2. Context-Aware Balancer
Simple fusion or weighted representations will not be able to adapt to dynamic changes in context. Therefore, the representations and are then integrated via a context-aware fusion gate that dynamically computes balancing weights. This gate learns to adjust the contribution of each branch based on the specific input instance, thereby mitigating the imbalance between noise and uncertainty. First, the two representations are element-wise add to form a unified context tensor . An instance-specific gating signal is then generated through a multi-step transformation:
where first aggregates spatial information to create a channel-wise descriptor. This descriptor is then passed through a bottleneck structure composed of two layers, with a ReLU non-linearity, and finally a activation to produce the balancing weight . The final prediction is obtained by a weighted combination of the two branches, where is broadcast appropriately:
In addition, we also use the residual connection technique to supplement the context tensor to promote information propagation. This adaptive gating mechanism allows the model to dynamically prioritize past versus future information based on the input context, effectively balancing their distinct type effects and producing a robust, fused result for prediction.
4. Experiments
@@
In this section, we conduct extensive experiments to investigate the following research questions:
@@
RQ1: Can ExoST outperform exogenous variable modeling methods in accuracy across various task scenarios? (Effectiveness)
RQ2: Can ExoST achieve consistent improvements on various types of existing spatio-temporal backbones? (Universality)
@@
RQ3: Can ExoST maintain its performance in scenarios where exogenous variables are missing or noisy? (Robustness)
@@
RQ4: How does ExoST work? Which components or strategies are crucial? Are these components or strategies sensitive to different desgin or hyper-parameters? (Mechanism)
@@
RQ5: How does ExoST compare in efficiency and parameter count to other methods? (Efficiency & Lightweight)
4.1. Experimental Setup
@@
Datasets. We conducted various experiments on two public real-world spatio-temporal datasets 111https://zenodo.org/records/7308425, which contains hourly measurements collected from monitoring stations in Madrid during the periods 1 January to 30 June 2019 and 1 January to 30 June 2022. Each dataset includes monitoring data from air quality, meteorological, and traffic systems. We consider four forecasting tasks: (i) AQI-19 and AQI-22: predict concentrations in air quality system using 2019 / 2022 data, with traffic and meteorological indicators as exogenous variables; (ii) Speed-19 and Intensity-22: predict traffic speed / intensity in the traffic system using 2019 / 2022 data, with other system indicators as exogenous variables. All dataset are chronologically split into training, validation, and test sets in a 7:2:1 ratio. In addition, it is worth noting that we also include the coding of dates as both past and future exogenous variables as input. More details on dataset, please refer to Appendix A.1.
@@
Baseline. For performance comparison, we categorize existing baselines into three groups: (i) time series models with exogenous variable modeling : TimeXer (Wang et al., 2024), TiDE (Das et al., 2023), and NBEATSx (Olivares et al., 2023); (ii) spatio-temporal models with exogenous variable integration: MAGCRN (Giganti et al., 2024); (iii) general exogenous variable modeling frameworks: ChronosX (Arango et al., 2025), which allows for the substitution of any backbone. Additionally, to evaluate ExoST’s universality, we extensively utilize various classic spatio-temporal backbones, including AGCRN (Bai et al., 2020), GWNet (Wu et al., 2019), GGNN (Satorras et al., 2022), GRUGCN (Gao and Ribeiro, 2022), STGCN (Yu et al., 2017), and DCRNN (Li et al., 2017). More details on methods, see Appendix A.2.
Protocol. By default, we use the previous 24 steps of historical data to predict data for the next 24 steps (i.e., 1-day), and evaluate using mean absolute error (MAE), root mean square error (RMSE), mean relative error (MRE), and mean absolute percentage error (MAPE). Moreover, to measure the long-term generalization of the model in real-world deployments, we also perform rolling evaluations over 2 and 3 days without retraining. For the baselines, we use the parameter settings recommended in their papers and fine-tune them. For ExoST, we set the number of gated experts to 4 and the hidden size to 64. To ensure fairness, each experiment is repeated five times, and results are reported as the mean standard deviation. Best results are highlighted in bold pink and the second-best in underlined blue. For more details on the protocol, see Appendix A.3.
4.2. Effectiveness Study (RQ1)
@@
Table LABEL:tab:rq1_overall_performance shows the performance comparison of our model with various state-of-the-art methods for time series forecasting or spatio-temporal forecasting that utilize exogenous variable modeling. It is worth noting that the first three time series forecasting methods have coupled model designs, while MAGCRN, ChronosX, and our method all use classical AGCRN(Bai et al., 2020) as the backbone network to ensure a fair comparison. Our observations are as follows: ❶ Time series forecasting methods generally achieve the worst results, which is consistent with common spatio-temporal forecasting benchmarks (Shao et al., 2024). This is because they generally do not consider spatial dependencies, resulting in inaccurate predictions. ❷ The ChronosX method, which utilizes a spatio-temporal backbone, despite achieving relative progress, often encounters non-convergence issues on some tasks during training (i.e., Speed-19 and Intensity-22). Despite repeated parameter tuning, this issue cannot be avoided. We attribute this to the design of the underlying Chronos-type (Ansari et al., [n. d.]) time series foundation model. ❸ MAGCRN often achieves suboptimal results because it is specifically designed for spatio-temporal prediction scenarios and still ignores the inconsistent and unbalanced effects of exogenous variables. ❹ Our ExoST achieves competitive performance across all metrics on the four tasks. By effectively selecting and balancing exogenous information, the ExoST can adapt to a variety of input sources, maintaining stable performance.
4.3. Universality Study (RQ2)
We further aim to demonstrate the universality of our ExoST’s “select, then balance” paradigm across various spatio-temporal backbones. Table LABEL:tab:rq1_3day illustrates the performance improvements of our proposed ExoST for 3-day predictions across four tasks and six models. The ✗ column reflects the the performance metrics of the backbone model alone on the target system, while the ✓ column displays the improved results achieved by our ExoST through exogenous variable modeling. The relative improvement is quantified in the column. Based on these results, we make the following observations: ❶ Our ExoST exhibits significant performance improvements across different architectures and datasets, attributed to an effective exogenous variable modeling paradigm. ❷ From the model perspective, the extent of performance improvement varies among different backbone networks, with AGCRN yielding the optimal overall enhancement; however, even the least performing Gwnet model achieved substantial performance gains. ❸ From the task perspective, air quality prediction consistently showed stable performance improvements, as exogenous information like traffic and meteorology significantly impacts it; conversely, traffic speed prediction saw only moderate improvements due to the limited influence of exogenous information such as air quality. We further supplement results for 1- and 2-day predictions in Appendix B.1 and additionally find that ❹ as the prediction horizon extends, our ExoST achieves more significant gains, which also indicates the strong generalization capability of our method.
4.4. Robustness Study (RQ3)
@@
In complex real-world scenarios, exogenous variable data in spatio-temporal systems may not always be collected reliably (for example, signal loss caused by sensor anomalies or maintenance), which can affect prediction accuracy and model robustness. To evaluate the robustness of the ExoST under such adverse conditions, we designed a systematic experiment to simulate real-world data corruption scenarios. Specifically, we randomly replaced exogenous variables with varying degrees of degradation (20%, 40%, 60%, and 80%) with zero values or randomly generated values from a standard normal distribution to evaluate the model’s ability to maintain predictive performance despite exogenous information corruption. The 3-day prediction experiment results shown in Table LABEL:tab:mask_3day reveal several interesting phenomena: ❶ As the proportion of missing data increases, model performance gradually decreases in most cases. However, the overall absolute performance drop is not significant, indicating that our framework can still maintain reasonable performance with limited auxiliary information from exogenous variables. ❷ For some tasks, appropriate missing signals can lead to improved performance, primarily because random data corruption can be viewed as a simple data augmentation strategy. By introducing different data perturbations, the model is forced to learn to predict under a variety of incomplete or noisy data conditions, thereby enhancing its generalization and robustness. We also provide a performance comparison for the 1-day and 2-day scenarios in Appendix B.2, observing similar results.
4.5. Mechanism Study (RQ4)
4.5.1. Ablation Study
@@
We first conducted an ablation study to systematically evaluate the impact of different components on ExoST from two perspectives: data input and model component.
@@
From the data perspective, we aimed to investigate the auxiliary role of different types of exogenous variables in the final prediction. Specifically, due to the universality of date information, here, we treated it as a separate category. As shown in Table LABEL:tab:performance_1day, we observed that ❶ future exogenous variables outperformed past, indicating that forward-looking information provides more direct predictive signals. However, combining both past and future yielded better performance than using either type alone, which suggests that the combination can provide complementary patterns. ❷ We also found that date exogenous variables performed poorly when used in isolation but led to improvements when combined with any other type. ❸ Best performance was achieved by combining all exogenous variables. This demonstrates that ExoST can effectively exploit different information, extracting and balancing contributions from different temporal perspectives and contextual information.
@@
From the model perspective, we aimed to investigate the auxiliary role of different module components in the final prediction. We constructed two variants by removing the MoE Selector and the Context-aware Balancer (achieved through direct element-wise fusion), respectively. As Figure 9 shows, ❶ w/ Selector variant significantly degrades performance in all metrics, demonstrating the critical importance of addressing inconsistent variable effects. Performance degradation indicates that without proper semantic alignment and adaptive selection mechanisms, the model struggles to disentangle heterogeneous exogenous variables from endogenous variables, leading to feature space competition and channel entanglement issues. ❷ w/ Balancer variant causes even more pronounced performance degradation than w/ Selector variant, highlighting that temporal distribution differences between past and future exogenous information pose a more fundamental challenge, requiring sophisticated balancing mechanisms to handle noise-contaminated past information and uncertainty-amplified future forecasts.
More experimental results are shown in Appendix B.3.1.
4.5.2. Parameter Study.
@@
We then conducted a parameter study to evaluate the impact of different parameter on ExoST. As Figure 5 shows, we can observe that ❶ The hidden size of 64 achieves optimal performance in all prediction horizons, with larger sizes leading to performance degradation due to overfitting. ❷ Expert number of 4 provides the best balance between model capacity and computational efficiency. Too few experts (one or two) fail to capture the diverse semantic patterns needed for disentangling inconsistent variable effects, while excessive experts (eight) lead to overselection where redundant experts compete for similar patterns, diminishing the effectiveness of the mixture-of-experts gating mechanism designed to mitigate channel entanglement. These observations validate the effectiveness of our lightweight design choices in the latent-space gated expert module.
4.5.3. Strategy Study.
@@
We also use several alternative strategies to test the rationality of our design. These include: (i) Shared Parameter: Two siamese branches share parameters to learn a unified spatio-temporal representation. (ii) Simple Weight: Adopts a predefined balancing weight () to fuse the outputs from two siamese branches. (iii) Learnable Weight: Dynamically adjusts the relative importance of past and future information through learnable parameters. (iv) Attention: Utilizes cross-attention mechanisms to dynamically balance the contributions from two siamese branches. As shown in Figure 6, we observe that: ❶ ExoST achieves competitive performance in other strategies, demonstrating the effectiveness of our design choice. ❷ Simple Weight shows significant performance degradation due to its static balancing strategy that does not adapt to varying input contexts. ❸ Attention and Learnable Weight strategies underperform while increasing computational complexity, as attention mechanisms introduce excessive parameter overhead and learnable weights lack the sophisticated context modeling capabilities needed to handle nuanced differences between exogenous variables of the past and future. More experimental results are shown in Appendix B.3.2.
4.6. Efficiency & Lightweight Study (RQ5)

### v2

@@ 4.2 Balance: Dual-Branch Adaptive Fusion
where encode high-level spatio-temporal features derived from past and future exogenous contexts, respectively, facilitating subsequent fusion.
Context-Aware Balancer. To address the limitations of static fusion strategies, we employ a dynamic context-aware mechanism to integrate and . We first construct a unified context tensor , from which an instance-specific balancing weight is derived via:
Here, aggregates spatial information to generate a channel descriptor, which passes through an MLP bottleneck with Sigmoid activation . The final prediction is synthesized using a residual connection to facilitate information propagation:
This adaptive calibration effectively prioritizes past versus future contexts based on instance-specific dynamics, yielding a robust and balanced prediction.
Analysis. Prior modeling of past and future exogenous variables typically relies on static fusion, employing either fixed linear weights or global affine transformations optimized over the training distribution (das2023long; giganti2024back; Arango2025ChronosX). In contrast, our context-aware mechanism utilizes input-dependent non-linear mappings to approximate arbitrary fusion functions, enabling instance-specific adaptation during inference (More formal analysis in Appendix A.2). Fig. 4 visualizes the weight distributions across inference samples in the AQI-19 experiment; unlike the scattered and irregular patterns observed in baselines, ExoST exhibits smoother and more uniform distributions, underscoring its superior dynamic adaptability. Empirically, we also validate the effectiveness of this design in Section. 5.5.
5 Experiments
@@
In this section, we conduct extensive experiments to investigate the following core research questions:
RQ1: (Universality) Can ExoST achieve consistent improvements on any existing ST backbones?
@@
RQ2: (Effectiveness) Can ExoST outperform existing other methods across various task scenarios?
@@
RQ3: (Robustness) Can ExoST maintain its performance when exogenous variables are missing or noisy?
@@
RQ4: (Mechanism & Rationality) How does ExoST work? Which components or strategies are crucial, and are they sensitive to different designs or hyper-parameters?
@@
RQ5: (Efficiency & Lightweight) How does ExoST compare in efficiency and parameter count to others?
5.1 Experimental Setup
@@
Datasets. We evaluate on several public real-world spatio-temporal datasets 111https://zenodo.org/records/7308425, containing hourly measurements from air quality, meteorological, and traffic monitoring stations in Madrid (1 January–30 June, 2019/2022). We study four forecasting tasks: (i) AQI-19 and AQI-22: predict in air quality domain concentrations with traffic and meteorological domain variables as exogenous inputs; (ii) Speed-19 and Intensity-22: predict traffic speed / intensity using other system indicators as exogenous variables. Datasets are chronologically split into train/val/test with ratio 7:2:1. Date encodings are included as past and future exogenous variables. More dataset details are provided in Appendix B.1.
@@
Baseline. We compared with three comparable types of baselines in the literature: (i) time-series models with exogenous modeling: TiDE (das2023long), NBEATSx (olivares2023neural), TimeXer (wang2024timexer), CrossLinear (zhou2025crosslinear), DAG (qiu2025dagdualcausalnetwork); (ii) spatio-temporal models with exogenous integration: MAGCRN (giganti2024back); (iii) general exogenous modeling framework (for foundation model): ChronosX (Arango2025ChronosX). To test ExoST’s universality, we further instantiate it with various classic spatio-temporal backbones, including AGCRN (Bai2020Adaptive), GWNet (wu2019graph), GGNN (Satorras2022Multivariate), GRUGCN (Gao2022On), STGCN (yu2017spatio), and DCRNN (li2017diffusion). More details on baseline methods and backbone networks, please refer to Appendix B.2.
Protocol. We use the past 24 steps to forecast the next 24 steps (i.e., 1-day) and report MAE, RMSE, MAPE, and MRE. To assess long-term robustness, we additionally perform 2- and 3-day rolling forecasts without retraining. Baseline models follow their recommended settings. For ExoST, we use 4 gated experts and hidden size 64. Each experiment is run five times; we report mean standard deviation. Best results appear in bold blue, second-best in underlined green. Full protocol details can be found in Appendix B.3.
5.2 Universality Study (RQ1)
We first evaluate the universality of the “select, then balance” paradigm in ExoST across diverse ST backbones. Figure 5 illustrates the absolute and average relative performance improvements of ExoST across four tasks (left) and six models (right) for 3-day forecasting. ❶ ExoST yields consistent performance boosts across varying architectures and datasets, validating the efficacy of our exogenous modeling paradigm. ❷ From the task perspective, air quality forecasting achieves substantial gains due to the strong influence of exogenous factors (e.g., traffic and meteorology), whereas traffic speed forecasting sees moderate improvements, reflecting a limited dependency on exogenous variables. ❸ From the model perspective, while gains vary across backbones, most exceed 20%; notably, even the vanilla GWNet achieves a significant improvement of nearly 10%. ❹ Supplementary results for 1-day and 2-day horizons (Tables C.2 and C.3 in Appendix C.1) further reveal that performance gains amplify as the forecasting horizon extends, underscoring the robust generalization capability of our approach.
| Datasets | AQI-19 | Speed-19 ∗ | AQI-22 | Intensity-22 | |||||||||
| Method | Metric | 1-day | 2-day | 3-day | 1-day | 2-day | 3-day | 1-day | 2-day | 3-day | 1-day | 2-day | 3-day |
| TiDE (das2023long) | MAE | ||||||||||||
| RMSE | |||||||||||||
| MAPE (%) | |||||||||||||
| MRE (%) | |||||||||||||
| NBEATSx (olivares2023neural) | MAE | ||||||||||||
| RMSE | |||||||||||||
| MAPE (%) | |||||||||||||
| MRE (%) | |||||||||||||
| TimeXer (wang2024timexer) | MAE | ||||||||||||
| RMSE | |||||||||||||
| MAPE (%) | |||||||||||||
| MRE (%) | |||||||||||||
| CrossLinear (zhou2025crosslinear) | MAE | ||||||||||||
| RMSE | |||||||||||||
| MAPE (%) | |||||||||||||
| MRE (%) | |||||||||||||
| DAG (qiu2025dagdualcausalnetwork) | MAE | ||||||||||||
| RMSE | |||||||||||||
| MAPE (%) | |||||||||||||
| MRE (%) | |||||||||||||
| MAGCRN (giganti2024back) | MAE | ||||||||||||
| RMSE | 108.66±17.87 | ||||||||||||
| MAPE (%) | 42.53±0.41 | 43.79±4.21 | 45.89±4.35 | 49.74±4.19 | |||||||||
| MRE (%) | |||||||||||||
| ChronosX (Arango2025ChronosX) | MAE | ||||||||||||
| RMSE | |||||||||||||
| MAPE (%) | |||||||||||||
| MRE (%) | |||||||||||||
| ExoST | MAE | 9.33±0.13 | 10.04±0.19 | 10.41±0.12 | 3.79±0.01 | 3.80±0.01 | 3.80±0.01 | 7.33±0.21 | 7.61±0.23 | 7.94±0.29 | 31.91±1.35 | 50.59±0.04 | 64.13±3.86 |
| RMSE | 14.24±0.18 | 15.12±0.26 | 15.51±0.16 | 31.73±0.02 | 31.78±0.02 | 31.85±0.02 | 11.14±0.21 | 11.52±0.49 | 11.79±0.51 | 56.49±2.54 | 89.10±7.36 | ||
| MAPE (%) | 38.57±0.32 | 44.65±0.08 | 9.30±0.03 | 9.32±0.03 | 9.31±0.03 | 17.63±0.97 | 26.68±1.69 | 34.52±2.90 | |||||
| MRE (%) | 36.09±0.49 | 38.84±0.75 | 40.34±0.47 | 9.88±0.03 | 9.89±0.03 | 9.90±0.03 | 39.23±1.11 | 40.74±1.24 | 42.50±1.57 | 11.57±0.49 | 18.31±1.14 | 23.16±1.40 | |
5.3 Effectiveness Study (RQ2)
@@
Table 2 presents a performance comparison of our ExoST against state-of-the-art time-series, spatio-temporal, and foundation models with exogenous variable modeling techniques. To ensure a fair comparison, we employ the standard AGCRN (Bai2020Adaptive) as the backbone for all methods utilizing a decoupled design. Our observations are as follows: ❶ Pure time-series methods (first 5) consistently yield the lowest performance, aligning with established ST benchmarks (shao2024exploring). This deficiency stems from their inability to model spatial dependencies, resulting in suboptimal accuracy. ❷ While ChronosX achieves relative gains by integrating the spatio-temporal backbone, it frequently suffers from non-convergence on specific tasks (e.g., Speed-19 and Intensity-22), despite extensive hyperparameter tuning. We attribute this instability to the inherent design of the underlying Chronos-type foundation models (ansarichronos). ❸ MAGCRN often produces suboptimal results as it is tailored for specific scenarios yet neglects the inconsistent and unbalanced effects of exogenous variables. ❹ Conversely, our ExoST achieves superior performance across all metrics on the four tasks. By effectively selecting and balancing exogenous information, ExoST adaptively handles diverse input sources to maintain robust performance.
5.4 Robustness Study (RQ3)
@@
In real-world scenarios, spatio-temporal systems frequently encounter unreliable exogenous data collection (e.g., signal loss from sensor anomalies), which compromises prediction accuracy and robustness. To assess the ExoST under such conditions, we designed a systematic experiment simulating data corruption. Specifically, we randomly replaced exogenous variables at varying degradation levels (20%, 40%, 60%, and 80%) with zeros or noise from a standard normal distribution . The 3-day prediction results in Table 3 reveal that: ❶ While performance gradually declines as the missing data ratio increases, the absolute drop remains minimal, indicating our framework maintains reasonable efficacy even with limited auxiliary information. ❷ Surprisingly, for certain tasks, appropriate levels of missing signals improve performance. This suggests that random corruption acts as a data augmentation strategy, forcing the model to predict under incomplete or noisy conditions, thereby enhancing generalization and robustness. Similar trends for 1-day and 2-day scenarios are reported in Appendix C.2.
| Datasets | AQI-19 | Speed-19∗ | Intensity-22 | ||||||||||
| Zero | 20% | 108.23±18.00 | 34.19±3.45 | ||||||||||
| 40% | |||||||||||||
| 60% | |||||||||||||
| 80% | |||||||||||||
| Random | 20% | ||||||||||||
| 40% | |||||||||||||
| 60% | |||||||||||||
| 80% | |||||||||||||
| No Masking | 10.41±0.12 | 15.51±0.16 | 42.62±0.52 | 38.84±0.75 | 3.80±0.01 | 31.85±0.02 | 9.31±0.03 | 9.90±0.03 | 64.13±3.86 | 23.16±1.40 | |||
5.5 Mechanism & Rationality Study (RQ4)
@@
Ablation Study. We conducted an ablation study to systematically evaluate the impact of different components on ExoST from both data and model perspectives.
| P | F | D | MAE | RMSE | MAPE (%) | MRE (%) |
| ✓ | ✓ | ✓ | 9.33±0.13 | 14.24±0.18 | 38.57±0.32 | 36.09±0.49 |
@@
From the data perspective, we investigated the auxiliary role of distinct exogenous variable types. Given the universality of date information, we treat it as a separate category. As shown in Table 6: ❶ Future exogenous variables outperform past variables, indicating that forward-looking information provides more direct predictive signals; however, combining both yields superior performance, suggesting they offer complementary patterns. ❷ Date exogenous variables perform poorly in isolation but drive improvements when combined with other types. ❸ The best performance is achieved by integrating all exogenous variables, demonstrating that ExoST effectively extracts and balances contributions from diverse temporal contexts.
@@
From the model perspective, we examined the contribution of specific modules by constructing variants w/o the Gated Expert Selector or the Context-aware Balancer (replaced by direct element-wise fusion). As Figure 6 illustrates: ❶ The w/o Selector variant suffers significant degradation across all metrics, proving the importance of addressing inconsistent variable effects. Without semantic alignment and adaptive selection, the model struggles to disentangle heterogeneous exogenous variables from endogenous ones, leading to feature space competition. ❷ The w/o Balancer variant exhibits even more pronounced degradation, highlighting that temporal distribution shifts between past and future exogenous information pose a fundamental challenge. Sophisticated balancing mechanisms are key to manage noise-contaminated data and uncertainty-laden future forecasts. More experimental results are shown in Appendix C.3.1.
@@
Parameter Study. We then performed a parameter sensitivity analysis to assess the impact of hyper-parameters on ExoST. As shown in Figure 8: ❶ A hidden size of 64 achieves optimal performance across all prediction horizons, with larger sizes causing degradation due to overfitting. ❷ Setting the number of experts to 4 provides the best trade-off between model capacity and computational efficiency. Fewer experts fail to capture the diverse semantic patterns required to disentangle inconsistent effects, while excessive experts lead to redundancy and overselection, diminishing the effectiveness of the mixture-of-experts gating mechanism. These findings validate the lightweight design choices of our latent-space gated expert module.
@@
Strategy Study. We compared our approach against several alternative fusion strategies to validate our design rationale:(i) Shared Parameter: Siamese branches sharing parameters to learn a unified representation.(ii) Simple Weight: A static balancing weight () fusing outputs.(iii) Learnable Weight: Dynamically adjusting importance via learnable parameters.(iv) Attention: Using cross-attention to balance contributions.As shown in Figure 8: ❶ ExoST outperforms all alternative strategies, confirming the effectiveness of our design.❷ Simple Weight degrades performance significantly due to its inability to adapt to varying input contexts.❸ Attention and Learnable Weight strategies underperform while increasing computational cost; attention mechanisms introduce excessive parameter overhead, and learnable weights lack the sophisticated context modeling required to handle the nuanced differences between past and future exogenous variables. More experimental results are shown in Appendix C.3.2.
5.6 Efficiency & Lightweight Study (RQ5)